OpenStudy (anonymous):

The circle (x−2)^2+(y−3)^2=4 can be drawn with parametric equations. Assume the circle is traced clockwise as the parameter increases. If x=2+2cos(t) then y =

OpenStudy (anonymous):

(2 + 2 cos (t) - 2)² + (y - 3)² = 4 4 cos² t + (y - 3)² = 4 Solve for y

OpenStudy (anonymous):

5-4cos(t) i got this but it wasn't correct still

OpenStudy (anonymous):

can u help?

OpenStudy (anonymous):

(y - 3)² = 4 - 4 cos² t (y - 3)² = 4(1 - cos² t) \[y - 3 = \pm \sqrt{4(1-\cos^2 t)}\] \[y = 3\pm 2\sqrt{(1-\cos^2 t)}\]

OpenStudy (anonymous):

awesome thank u for ur help

OpenStudy (anonymous):

np